you purchase 3 squash racquets, 7 pairs of socks and 4 head-bands from a sports shop. You hand the cashier R2500 and receive R250 in change. A squash racquet costs R510 more than a pair of socks, which in turn is half the price of a head-band
1
Expert's answer
2020-06-12T10:23:24-0400

Let "x=" the cost of one pair of socks. Then the cost of one squash racquet is "x+510," the cost of one head-band is "2x"

You purchase 3 squash racquets, 7 pairs of socks and 4 head-bands from a sports shop. You hand the cashier R2500 and receive R250 in change.


"2500-(3(x+510)+7x+4(2x))=250"

Solve for "x"


"2500-3x-1530-7x-8x=250"

"18x=720"

"x=40"

Cost of one squash racquet "=40+510=550"

Cost of one pair of socks "=40"

Cost of one head-band "=2(40)=80"


Cost of one squash racquet is R550.

Cost of one pair of socks is R40.

Cost of one head-band is R80.